A 16 foot ladder reaches a window 14 feet abouve the ground. How far from the wall is the base of the ladder. Round your answer to the nearest whole number.
Answer by checkley79(3341)

A^2+B^2=C^2
14^2+B^2=16^2
B^2=16^2-14^2
B^2=256-196
B^2=60
B=SQRT60
B=8 ANS.